# Entropy Stability and Spectral Concentration under Convex Block Constraints

### v1

@@ Abstract
We investigate entropy minimization problems for quantum states subject to convex block-diagonal constraints. Our principal result is a quantitative stability theorem: if a state has entropy within $\varepsilon$ of the minimum possible value under a fixed block constraint, then it must lie within $O(\sqrt{\varepsilon})$ trace distance of the manifold of entropy minimizers. We prove that this stability rate is optimal.
The analysis is entirely finite-dimensional and relies on a precise decomposition of entropy into classical and internal components, together with sharp relative entropy inequalities. As a concrete application, we study self-adjoint operators with finite spectral decompositions whose eigenspaces induce natural block constraints. In this setting, the stability theorem yields quantitative non-concentration bounds for the associated spectral measures.
The framework is abstract and independent of arithmetic or dynamical input. It provides a general stability principle for entropy minimizers under structured convex spectral constraints.
@@ 1 Introduction
In this paper we study the stability of entropy minimizers under a natural class of convex constraints arising from orthogonal decompositions of a Hilbert space. While the structure of exact entropy minimizers is well understood in many constrained settings, considerably less is known about how close a state with nearly minimal entropy must be to the set of minimizers, measured in a natural metric such as trace norm.
The principal goal of this work is to establish a sharp quantitative stability theorem of the following form: if a quantum state satisfies a fixed block-diagonal constraint and its entropy exceeds the minimum allowed by the constraint by at most $\varepsilon$, then the state lies within $O(\sqrt{\varepsilon})$ in trace norm of the manifold of entropy minimizers. We prove that this exponent is optimal in general and cannot be improved under the given hypotheses.
@@ 1.1 Background and motivation
In practice, however, one rarely encounters exact minimizers. Instead, one typically works with states whose entropy is close—but not equal—to the minimum allowed by the constraints. This leads to the following natural stability question:
@@
This is a stability problem rather than a minimization problem: it asks for explicit control of distance to the minimizing set in terms of the entropy gap, rather than identification of the minimizers themselves.
@@ 1.2 Block-diagonal constraints
For block-diagonal states, the entropy admits a canonical decomposition into a classical (Shannon) contribution arising from block weights and a quantum contribution arising from the internal states within each block. This structural decomposition enables a precise analysis of entropy minimization and stability.
@@ 1.3 Main result (informal)
Among all block-diagonal states with $r$ blocks, the minimal entropy equals $\log r$ and is achieved precisely by uniform mixtures of pure states, one from each block. This description follows directly from the entropy decomposition and the strict concavity of Shannon entropy.
Our main result shows that this characterization is quantitatively stable:
@@ 1.4 Relation to existing work
Stability phenomena for entropy and related functionals have been studied in several contexts, including Pinsker-type inequalities for relative entropy, logarithmic Sobolev inequalities, and stability results in convex optimization. These results typically control relative entropy with respect to a fixed reference state or quantify entropy decay along dynamical flows.
@@ 1.5 Structure of the paper
Section 2 introduces notation and recalls basic entropy inequalities. Section 3 analyzes the geometry of block-diagonal convex constraint sets and characterizes entropy minimizers. Section 4 establishes the main quantitative stability theorem with complete proof. Section 5 proves optimality of the stability exponent through explicit classical and quantum examples. Section 6 discusses interpretations and consequences of the stability theorem. The appendix collects standard entropy facts used throughout the paper.
@@ 2 Preliminaries and Notation
Throughout the paper, $\mathcal{H}$ denotes a finite-dimensional complex Hilbert space of dimension $d\geq 1$. All operators are linear and act on $\mathcal{H}$ unless otherwise stated. Finite dimensionality is assumed throughout and is used only to guarantee compactness and norm equivalence; no dimension-dependent constants appear in the main results.
@@ 2.2 Von Neumann entropy and relative entropy
The von Neumann entropy of a state $\rho\in\mathcal{S}(\mathcal{H})$ is defined by
@@
where the logarithm is taken in the functional calculus sense, with the convention $0\log 0=0$.
The entropy satisfies the basic bounds
@@
$D(\rho\|\sigma):=\operatorname{Tr}\!\big(\rho(\log\rho-\log\sigma)\big).$
If $\operatorname{supp}(\rho)\nsubseteq\operatorname{supp}(\sigma)$, we set $D(\rho\|\sigma)=+\infty$.
Relative entropy is jointly convex, nonnegative, and vanishes if and only if $\rho=\sigma$. It will be used exclusively as a technical tool in stability arguments.
@@ 2.3 Entropy inequalities
We recall several standard inequalities that will be used repeatedly.
@@
where $H(p):=-\sum_{i}p_{i}\log p_{i}$ is the Shannon entropy.
Proof.
Since the blocks act on orthogonal subspaces, the spectrum of $\rho$ is the union of the spectra of the operators $p_{i}\rho_{i}$. The identity follows by direct calculation. ∎
@@
Remark 2.3.
Sharper constants are known, but the stated form suffices for all arguments in this paper.
@@ 2.4 Purity and effective dimension
Proof.
This follows from Jensen’s inequality applied to the convex function $x\mapsto-\log x$ and the spectral decomposition of $\rho$. ∎
@@ 2.5 Constrained entropy minimization
A state $\rho\in\mathcal{C}$ is called an entropy minimizer if $S(\rho)=S_{\min}(\mathcal{C})$. The main problem studied in this paper is the stability of entropy minimizers: given $\rho\in\mathcal{C}$ with $S(\rho)-S_{\min}(\mathcal{C})$ small, how close must $\rho$ be (in trace norm) to the set of minimizers?
@@ 3.3 Block-convex constraint sets
We now define the class of constraint sets studied in this paper.
@@
Since $\Pi$ and each $\mathcal{C}_{i}$ are convex, we have $\lambda p+(1-\lambda)q\in\Pi$ and $\tilde{\rho}_{i}\in\mathcal{C}_{i}$. Compactness follows from finite dimensionality. ∎
@@ 3.4 Entropy minimizers: marginal structure
The Shannon entropy $H$ is strictly concave on every face of $\Delta_{r}$. The second term is affine in $p$. If $p$ were not extreme in $\Pi$, then $p$ would be a nontrivial convex combination of distinct points in $\Pi$, strictly lowering $H(p)$ at one endpoint and contradicting minimality. ∎
@@ 3.5 Entropy minimizers: conditional structure
Proof.
Fix $i$ with $p_{i}>0$. Suppose $\rho_{i}$ is not an entropy minimizer in $\mathcal{C}_{i}$. Then there exists $\sigma_{i}\in\mathcal{C}_{i}$ with $S(\sigma_{i})<S(\rho_{i})$. Replacing $\rho_{i}$ by $\sigma_{i}$ while keeping all other blocks unchanged yields a state in $\mathcal{C}$ with strictly smaller entropy, contradicting minimality. ∎
@@ 3.7 Compactness of the minimizer set
Proof.
The set $\mathcal{C}$ is compact, and entropy is lower semicontinuous. Hence the set of minimizers is closed and therefore compact. ∎
@@ 3.8 Preliminary continuity estimate
Proof.
This follows from compactness of $\mathcal{M}$ and lower semicontinuity of entropy. ∎
@@ 4 A Quantitative Stability Theorem
This section establishes the main result of the paper: entropy minimizers under block-convex constraints are quantitatively rigid. We show that entropy excess controls the squared trace-norm distance to the minimizer set, with constants depending only on the geometry of the constraint.
@@ 4.2 Statement of the stability theorem
There exists a constant $C>0$, depending only on the geometry of the marginal polytope $\Pi$ and the conditional constraint sets $\{\mathcal{C}_{i}\}$, such that for all $\rho\in\mathcal{C}$,
@@
Remark 4.2.
The constant $C$ in Theorem 4.1 depends only on the geometry of the constraint, namely on the number of blocks $r$, the combinatorial geometry of the marginal polytope $\Pi$, and the convex sets $\{\mathcal{C}_{i}\}$. In particular, $C$ is independent of the internal block dimensions $d_{i}=\dim(\mathcal{H}_{i})$ and of the ambient Hilbert space dimension.
@@ 4.4 Marginal entropy stability
Let $\Pi\subset\Delta_{r}$ be a compact convex polytope and let $q\in\mathrm{ext}(\Pi)$. Then there exists a constant $c_{1}>0$, depending only on $\Pi$, such that for all $p\in\Pi$,

### v2

# Quadratic Stability of Entropy Minimizers under Block-Separable Convex Constraints

@@ Abstract
We study quadratic stability of entropy minimization under convex constraints with block-separable structure. For constraint sets of the form $\mathcal{C}=\{\bigoplus_{i}p_{i}\rho_{i}:p\in\Pi,\ \rho_{i}\in\mathcal{C}_{i}\}$, we establish explicit stability estimates showing that, under a confining (fixed-support) hypothesis, the entropy gap controls the squared trace-norm distance to the set of entropy minimizers, with constants determined by the geometry of the constraint.
The stability constant admits a natural decomposition into marginal and conditional components: the marginal contribution is governed by the curvature of Shannon entropy on the marginal polytope $\Pi$ at extreme points, while the conditional contribution is governed by the curvature of von Neumann entropy at entropy-minimizing states within each block.
We prove that the quadratic exponent is optimal by constructing explicit examples for which no linear stability bound holds uniformly. Exploiting the block-separable structure, the analysis reduces the global stability problem to independent marginal and conditional subproblems, yielding a geometric characterization of the stability constants in terms of constraint curvature. This stability phenomenon cannot be derived from Pinsker–type inequalities or standard entropy continuity bounds, since no reference state is fixed and the minimizer emerges intrinsically from the constraint geometry.
keywords:
convex optimization, entropy minimization, quadratic stability, block-separable constraints, error bounds, explicit constants
{AMS}
90C25, 94A17, 49J52, 49J45
@@ 1 Introduction
We study the stability of entropy minimizers under a natural class of convex constraints arising from orthogonal decompositions of a Hilbert space. While the structure of exact entropy minimizers is well understood in many constrained settings, considerably less is known about how close a state with nearly minimal entropy must be to the set of minimizers, measured in a natural metric such as the trace norm.
The principal goal of this work is to establish a sharp quantitative stability theorem of the following form: if a quantum state satisfies a fixed block-diagonal constraint and its entropy exceeds the minimum allowed by the constraint by at most $\varepsilon$, then the state lies within $O(\sqrt{\varepsilon})$ in trace norm of the set of entropy minimizers. We prove that this exponent is optimal in general and cannot be improved under the given hypotheses. Purely quadratic entropy–distance stability holds under fixed-support (confining) constraints; see Appendix A for a precise geometric formulation.
@@ 1.1 Background and motivation
In practice, however, one rarely encounters exact minimizers. Instead, one typically works with states whose entropy is close, but not equal, to the minimum allowed by the constraints. This leads to the following natural stability question:
@@
This is a stability problem rather than a minimization problem: it asks for explicit control of distance to the minimizing set in terms of the entropy gap, rather than identification of the minimizers themselves. From an optimization viewpoint, the stability problem considered here is closely related to quadratic growth and error-bound properties for convex functionals. The result shows that, under a confining (fixed-support) constraint, the entropy functional exhibits quadratic growth relative to the set of minimizers when measured in trace norm. In this sense, the stability estimate can be interpreted as a second-order error bound for entropy minimization on a structured convex set, with constants determined explicitly by the geometry of the constraint.
@@ 1.2 Block-diagonal constraints
For block-diagonal states, the entropy admits a canonical decomposition into a classical (Shannon) contribution arising from block weights and a quantum contribution arising from the internal states within each block. This structural decomposition enables a precise analysis of entropy minimization and stability. Although the presentation uses quantum states for concreteness, no genuinely quantum phenomenon is invoked; the analysis applies verbatim to classical probability simplices and to finite-dimensional noncommutative state spaces.
The block structure considered here is not a technical artefact, but a concrete realisation of stratified convex geometry, which is the true object of the analysis.
@@ 1.3 Main result (informal)
Throughout this work, quadratic entropy–distance stability is established exclusively within the confining (fixed-support) regime. Boundary-transversal regimes, in which admissible perturbations explore directions transversal to the support of an entropy minimizer, exhibit a fundamentally different, non-quadratic behavior and are treated separately.
For a canonical class of block-diagonal constraints (see Section 6), the entropy minimization problem admits explicit solutions: in particular, the minimal entropy equals $\log r$ and is achieved by uniform mixtures of pure states, one from each block.
Here, dimension-independent means independent of the ambient Hilbert space dimension $\dim(\mathcal{H})$. The associated constants may depend on structural parameters fixed by the constraint geometry (e.g., the maximal active block dimension or minimal marginal weights), which are part of the model and remain uniform throughout the admissible set.
Our main result shows that this characterisation is quantitatively stable:
@@
1.4 Related Work and Contribution
Our stability result is related to several classical inequalities in information theory and convex optimization.
Pinsker’s inequality [Pinsker1964, Csiszar1967] provides a fundamental bound relating relative entropy to trace distance:
(1) $D(\rho\|\sigma)\geq\frac{1}{2}\|\rho-\sigma\|_{1}^{2}.$
On fixed-support strata where $\mathrm{supp}(\rho)\subseteq\mathrm{supp}(\sigma)$, the entropy gap is locally equivalent to the relative entropy $D(\rho\|\sigma)$ up to second order. In particular, Pinsker–type inequalities imply quadratic control of the trace distance in a neighbourhood of $\sigma$. In this sense, Theorem 4.1 is consistent with Pinsker’s inequality when a fixed reference state is present. The relationship between entropy differences and relative entropy has been further studied by Reeb [Reeb2015].
Classical continuity and metric inequalities for entropy include the works of Pinsker [Pinsker1964, Csiszar1967], Fannes [Fannes1973], Audenaert [Audenaert2007], Winter [Winter2016], and related developments in quantum information theory.
Entropy continuity has been extensively analyzed in the quantum information literature. Fannes [Fannes1973] established the first continuity bound for von Neumann entropy. Audenaert [Audenaert2007] provided sharp continuity estimates, and Winter [Winter2016] proved tight uniform continuity bounds valid across the entire state space. These results give global moduli of continuity for entropy as a function of states.
Our contribution relative to this prior work is threefold:
  1. 1.
    Geometric characterization of the quadratic regime: We provide a precise definition of the confining (fixed-support) regime (Definition 3.3) under which quadratic stability holds. This regime is characterized geometrically in terms of tangency to support strata and nondegeneracy of constrained entropy minimizers.
  2. 2.
    Explicit stability constants from constraint geometry: For block-separable constraint sets, we derive explicit formulas for the stability constant $C$ in Theorem 4.1. The constant decomposes into:
    • A marginal component $c_{1}$ determined by the Hessian of Shannon entropy $H(p)=-\sum_{i}p_{i}\log p_{i}$ restricted to faces of the marginal polytope $\Pi$,
    • A conditional component $c_{2}$ determined by the Hessian of von Neumann entropy $S(\rho_{i})$ at conditional entropy minimizers within each block $\mathcal{C}_{i}$.
    This geometric dependence on constraint curvature is made fully explicit in our analysis.
  3. 3.
    Optimality of the quadratic exponent: In Section 5, we prove that the exponent $2$ in the stability bound $S(\rho)-S_{\min}\geq C\|\rho-\sigma\|_{1}^{2}$ cannot be improved. We construct explicit families of states (both classical and quantum) for which the entropy gap scales precisely as $\varepsilon^{2}$ where $\varepsilon=\|\rho-\sigma\|_{1}$, precluding any uniform bound with exponent $\alpha<2$.
The block-separable structure studied here arises naturally in quantum systems with superselection rules, symmetry-reduced problems, and measurement-induced decoherence. Our analysis shows that this structure enables a complete decomposition of the stability problem into independent marginal and conditional subproblems.
Unlike Pinsker–type inequalities or entropy continuity bounds, our stability estimate does not compare two fixed states, nor does it rely on a reference distribution specified a priori. Instead, it controls the distance to an entropy minimizer that is defined intrinsically by the geometry of the constraint set itself and emerges only after solving the constrained variational problem.
@@ 1.5 Relation to existing work
Stability phenomena for entropy and related functionals have been studied in several contexts, including Pinsker–type inequalities for relative entropy, logarithmic Sobolev inequalities, and stability results in convex optimisation. These results typically control relative entropy with respect to a fixed reference state or quantify entropy decay along dynamical flows.
@@
Optimality
The quadratic entropy–distance exponent established in this work is optimal. Even in purely classical (commutative) settings, no uniform linear or subquadratic stability bound can hold in a neighbourhood of entropy minimizers. This limitation is intrinsic and is demonstrated by explicit counterexamples in Section 5.
1.6 Operational significance and scope
Beyond the abstract variational analysis, it is important to indicate how the geometric hypotheses underlying the stability theorem manifest in concrete settings. Section 8 therefore presents a minimal finite-dimensional model, motivated by leakage phenomena in quantum information processing, which illustrates the practical relevance of the confining versus boundary–transversal distinction introduced above.
The purpose of this example is not to derive additional theoretical results, but to show that entropy-based certification is effective precisely in the confining (fixed-support) regime and necessarily fails once transversal leakage directions are permitted.
@@ 1.7 Structure of the paper
Section 2 introduces notation and recalls basic entropy inequalities. Section 3 analyses the geometry of block-diagonal convex constraint sets and characterises entropy minimizers. Section 4 establishes the main quantitative stability theorem with a complete proof. Section 5 proves optimality of the stability exponent through explicit classical and quantum examples. Section 6 discusses interpretations and consequences of the stability theorem. Section 7 relates the results to classical entropy inequalities. Section 8 provides computational guidance, showing how the marginal and conditional stability constants can be computed from Hessian eigenvalues for specific constraint geometries. The appendices collect standard entropy facts and technical arguments used throughout the paper.
@@ 2 Preliminaries and Notation
Throughout, $\mathcal{H}$ denotes a finite-dimensional complex Hilbert space of dimension $d\geq 1$. All operators are linear and act on $\mathcal{H}$ unless otherwise stated. Finite-dimensionality is assumed solely to ensure compactness and equivalence of norms; no dimension-dependent constants appear in the main results.
Finite dimensionality is assumed throughout and is essential for compactness and norm equivalence.
@@ 2.2 Von Neumann entropy and relative entropy
The von Neumann entropy [vonNeumann1927, NielsenChuang] of a state $\rho\in\mathcal{S}(\mathcal{H})$ is defined by
@@
where the logarithm is taken in the functional calculus sense, with the convention $0\log 0=0$. It satisfies the bounds
@@
$D(\rho\|\sigma):=\operatorname{Tr}\!\big(\rho(\log\rho-\log\sigma)\big),$
and is set to $\infty$ otherwise. Relative entropy is nonnegative, jointly convex, and vanishes if and only if $\rho=\sigma$. It will be used only as a technical tool in the stability analysis.
@@ 2.3 Entropy inequalities
We recall a basic structural identity for block-diagonal states.
@@
where $H(p):=-\sum_{i}p_{i}\log p_{i}$ denotes the Shannon entropy.
Proof 2.2.
Since the blocks act on orthogonal subspaces, the spectrum of $\rho$ is the union of the spectra of the operators $p_{i}\rho_{i}$. The identity follows directly.
@@ 2.4 Purity and effective dimension
Proof 2.5.
This follows from Jensen’s inequality applied to the convex function $x\mapsto-\log x$ and the spectral decomposition of $\rho$.
@@ 2.5 Constrained entropy minimization
A state $\rho\in\mathcal{C}$ is called an entropy minimizer if $S(\rho)=S_{\min}(\mathcal{C})$. The central problem addressed here is the stability of entropy minimizers: given $\rho\in\mathcal{C}$ with $S(\rho)-S_{\min}(\mathcal{C})$ small, how close must $\rho$ be (in trace norm) to the set of minimizers?
@@ 3.3 Block-convex constraint sets
We now define the class of constraint sets considered here.
@@
Since $\Pi$ and each $\mathcal{C}_{i}$ are convex, we have $\lambda p+(1-\lambda)q\in\Pi$ and $\tilde{\rho}_{i}\in\mathcal{C}_{i}$. Compactness follows from finite-dimensionality.
@@ 3.4 Entropy minimizers: marginal structure
The Shannon entropy $H$ is strictly concave on every face of $\Delta_{r}$. The second term is affine in $p$. If $p$ were not extreme in $\Pi$, then $p$ would be a nontrivial convex combination of distinct points in $\Pi$, strictly lowering $H(p)$ at one endpoint and contradicting minimality.
@@ 3.5 Entropy minimizers: conditional structure
Proof 3.9.
Fix $i$ with $p_{i}>0$. Suppose $\rho_{i}$ is not an entropy minimizer in $\mathcal{C}_{i}$. Then there exists $\sigma_{i}\in\mathcal{C}_{i}$ with $S(\sigma_{i})<S(\rho_{i})$. Replacing $\rho_{i}$ by $\sigma_{i}$ while keeping all other blocks unchanged yields a state in $\mathcal{C}$ with strictly smaller entropy, contradicting minimality.
@@ 3.7 Compactness of the minimizer set
Proof 3.13.
The set $\mathcal{C}$ is compact, and entropy is lower semicontinuous. Hence the set of minimizers is closed and therefore compact.
@@ 3.8 Preliminary continuity estimate
Proof 3.15.
This follows from compactness of $\mathcal{M}$ and lower semicontinuity of entropy.
@@ 4 A Quantitative Stability Theorem
This section establishes the main result: entropy minimizers under block-convex constraints are quantitatively rigid. We show that entropy excess controls the squared trace-norm distance to the minimizer set, with constants depending only on the geometry of the constraint. Before proceeding to the local second-order analysis, we note that by continuity of the entropy functional on the compact constraint set and compactness of the set of entropy minimizers, there exists $\varepsilon_{0}>0$ such that any $\rho\in\mathcal{C}$ satisfying $S(\rho)-S_{\min}\leq\varepsilon_{0}$ lies in a neighbourhood of some entropy minimizer $\sigma\in\mathcal{M}$ where the fixed-support hypothesis holds and the local Taylor expansion applies.
@@ 4.1 Setup and notation
Throughout, the marginal polytope $\Pi$ is assumed to be a compact polytope with finitely many faces.
@@ 4.2 Statement of the stability theorem
Throughout, the confining (fixed-support) hypothesis is understood uniformly in a neighborhood of the minimizer set $\mathcal{M}$.
@@
Assume that the marginal minimizer satisfies Assumption 1 and that all entropy minimizers $\sigma\in\mathcal{M}$ satisfy the confining (fixed-support) hypothesis of Lemma 4.7.
Let $\mathcal{C}\subset\mathcal{S}(\mathcal{H})$ be a block–convex constraint set in the sense of Definition 3.3, induced by a fixed block decomposition of $\mathcal{H}$, and let $\mathcal{M}$ denote the set of entropy minimizers on $\mathcal{C}$.
Then there exists a constant $C=C(\Pi,\{\mathcal{C}_{i}\})>0$ such that for all $\rho\in\mathcal{C}$,
@@
Remark 4.2 (Scope and interpretation).
The theorem establishes quadratic entropy stability in the fixed-support (confining) regime. When constrained entropy minimizers lie on the boundary of state space and admissible perturbations explore directions transversal to the support, the quadratic stability mechanism used here necessarily fails or requires modification. Such boundary–transversal regimes are not addressed in the present work.
From an optimization viewpoint, Theorem 4.1 establishes a quadratic growth (error bound) property for the entropy functional restricted to a structured convex constraint set. The stability constant $C=C(\Pi,\{\mathcal{C}_{i}\})$ depends only on the geometry of the marginal polytope and the conditional constraint sets; an explicit form is discussed in Section 4.8.
Dependence of constants
The constant $C$ may depend on $d_{\max}:=\max_{i}\dim(H_{i})$ (or, equivalently, on the maximal active support rank imposed by the constraint geometry), but is independent of the ambient Hilbert space dimension once the block structure and confinement parameters are fixed.
@@ 4.4 Marginal entropy stability
Assumption 1 (Marginal confinement / fixed support in $\Pi$).
Let $q\in\operatorname{ext}(\Pi)$ be an entropy-minimizing marginal.
There exist an index set $I\subset\{1,\dots,r\}$ and a constant $\underline{q}>0$ such that:
  1. 1.
    $\operatorname{supp}(q)=I$ and $q_{i}\geq\underline{q}$ for all $i\in I$;
  2. 2.
    all admissible marginals $p\in\Pi$ in a neighborhood of $q$ satisfy $\operatorname{supp}(p)=I$ (equivalently, $\Pi$ is locally contained in the face $\Delta_{I}$).
@@
Assume Assumption 1. Let $\Pi\subset\Delta_{r}$ be a compact convex polytope and let $q\in\mathrm{ext}(\Pi)$. Then there exists a constant $c_{1}>0$, depending only on $\Pi$, such that for all $p\in\Pi$,
@@
Remark 4.5 (Local-to-global reduction for the marginal term).
The estimate in Lemma 4.3 is only used in the neighbourhood of the entropy-minimising extreme marginal $q$ selected in Section 4.3. If $\|p-q\|_{1}$ is bounded from below by a fixed constant, then H(p)H(q)H(p)-H(q) admits a strictly positive lower bound by compactness of $\Pi$, so the final stability inequality remains global after possibly decreasing the constant. Thus no loss of generality is incurred by interpreting Lemma 4.3 as a local rigidity statement around $q$.